# Towards deterministic non-Gaussianity on a chip

### v1

@@ Appendix D Supplemental material 2: Tensor and nonlinear parameter
$\gamma_{\mathrm{NL}}^{(1)}$ $\equiv\frac{3\sqrt{\omega_{P_{1}}\omega_{P_{1}}}}{4\epsilon_{0}v_{P_{1}}v_{P_{1}}\mathcal{L}^{(1)}}\int\mathrm{d}\bm{r}\Gamma_{ijkl}^{(3)}(\bm{r})\left[{\mathsf{d}}^{(1),i}_{S_{1}}(\bm{r}){\mathsf{d}}^{(1),j}_{I}(\bm{r})\right]^{*}{\mathsf{d}}^{(1),k}_{P_{1}}(\bm{r}){\mathsf{d}}^{(1),l}_{P_{1}}(\bm{r})e^{i(\kappa^{(1)}_{P_{1}}+\kappa^{(1)}_{P_{1}}-\kappa^{(1)}_{I}-\kappa^{(1)}_{S_{1}})\zeta},$ (50)
@@
$\gamma_{\mathrm{NL}}^{(2)}$ $\equiv\frac{3\sqrt{\omega_{I}\omega_{P_{2}}}}{4\epsilon_{0}v_{I}v_{P_{2}}\mathcal{L}^{(2)}}\int\mathrm{d}\bm{r}\Gamma_{ijkl}^{(3)}(\bm{r})\left[{\mathsf{d}}^{(2),i}_{S_{2}}(\bm{r}){\mathsf{d}}^{(2),j}_{S_{3}}(\bm{r})\right]^{*}{\mathsf{d}}^{(2),k}_{I}(\bm{r}){\mathsf{d}}^{(2),l}_{P_{2}}(\bm{r})e^{i(\kappa^{(2)}_{I}+\kappa^{(2)}_{P_{2}}-\kappa^{(2)}_{S_{3}}-\kappa^{(2)}_{S_{2}})\zeta}.$ (51)
To compute the nonlinear parameters, we relate the $\Gamma^{(3)}$ tensor to the nonlinear susceptibility tensor $\chi^{(3)}$
$\Gamma^{(3)}_{ijkl}(\bm{r})=\frac{\chi^{(3)}_{ijkl}(\bm{r})}{\epsilon_{0}^{2}\varepsilon_{1}^{4}(\bm{r})},$ (52)
@@
where in identifying the effective nonlinear coefficient we neglect any frequency dependence on the dielectric constants, since we consider frequencies in the same range [24]. We can relate the displacement fields to the electric fields
$\bm{\mathsf{d}}^{(\ell)}_{u}(\bm{r})=\epsilon_{0}\varepsilon_{1}(\bm{r})\bm{\mathsf{e}}^{(\ell)}_{u}(\bm{r}_{\perp},\zeta),$ (53)
and we can write
@@
$\begin{split}\gamma_{\mathrm{NL}}^{(\ell)}&=\frac{3\bar{\chi}^{(3)}\epsilon_{0}\sqrt{\omega_{{u_{3}}}\omega_{{u_{4}}}}}{4v_{{u_{3}}}v_{{u_{4}}}\mathcal{L}^{(\ell)}}\int\mathrm{d}\bm{r}_{\perp}\mathrm{d}\zeta\frac{\chi^{(3)}_{ijkl}(\bm{r})}{\bar{\chi}^{(3)}}\left[{\mathsf{e}}^{(\ell),i}_{{u_{1}}}(\bm{r}_{\perp}){\mathsf{e}}^{(\ell),j}_{{u_{2}}}(\bm{r}_{\perp})\right]^{*}{\mathsf{e}}^{(\ell),k}_{{u_{3}}}(\bm{r}_{\perp}){\mathsf{e}}^{(\ell),l}_{{u_{4}}}(\bm{r}_{\perp})e^{i(\kappa^{(\ell)}_{{u_{3}}}+\kappa^{(\ell)}_{{u_{4}}}-\kappa^{(\ell)}_{{u_{2}}}-\kappa^{(\ell)}_{{u_{1}}})\zeta},\end{split}$ (54)
where $\bm{{\mathsf{e}}}^{(\ell)}_{u}(\bm{r}_{\perp})$ can be extracted from an eigenmode solver for the specific structure and the integral over $\bm{r}_{\perp}$ can be evaluated numerically. We have used the fact that in the ring frame we can write the mode profiles as $\bm{\mathsf{d}}^{(\ell)}_{u}(\bm{r}_{\perp})$, since we assume the spatial profiles are independent of $\zeta$. For phase matched wavenumbers $\kappa^{(\ell)}_{{u_{3}}}+\kappa^{(\ell)}_{{u_{4}}}-\kappa^{(\ell)}_{{u_{2}}}-\kappa^{(\ell)}_{{u_{1}}}=0$, and we have
@@
$\begin{split}\gamma_{\mathrm{NL}}^{(\ell)}&=\frac{3\bar{\chi}^{(3)}\epsilon_{0}\sqrt{\omega_{{u_{3}}}\omega_{{u_{4}}}}}{4v_{{u_{3}}}v_{{u_{4}}}}\int\mathrm{d}\bm{r}_{\perp}\frac{\chi^{(3)}_{ijkl}(\bm{r})}{\bar{\chi}^{(3)}}\left[{\mathsf{e}}^{(\ell),i}_{{u_{1}}}(\bm{r}_{\perp}){\mathsf{e}}^{(\ell),j}_{{u_{2}}}(\bm{r}_{\perp})\right]^{*}{\mathsf{e}}^{(\ell),k}_{{u_{3}}}(\bm{r}_{\perp}){\mathsf{e}}^{(\ell),l}_{{u_{4}}}(\bm{r}_{\perp}).\end{split}$ (55)

### v2

@@ Appendix C Nonlinear parameters and tensors
$\gamma_{\mathrm{NL}}^{(1)}$ $\equiv\frac{3\sqrt{\omega_{P_{1}}\omega_{P_{1}}}}{4\epsilon_{0}v_{P_{1}}v_{P_{1}}\mathcal{L}^{(1)}}\int\mathrm{d}\boldsymbol{r}\Gamma_{ijkl}^{(3)}(\boldsymbol{r})\left[{\mathsf{d}}^{(1),i}_{S_{1}}(\boldsymbol{r}){\mathsf{d}}^{(1),j}_{I}(\boldsymbol{r})\right]^{*}{\mathsf{d}}^{(1),k}_{P_{1}}(\boldsymbol{r}){\mathsf{d}}^{(1),l}_{P_{1}}(\boldsymbol{r})e^{i(\kappa^{(1)}_{P_{1}}+\kappa^{(1)}_{P_{1}}-\kappa^{(1)}_{I}-\kappa^{(1)}_{S_{1}})\zeta},$ (38)
@@
$\gamma_{\mathrm{NL}}^{(2)}$ $\equiv\frac{3\sqrt{\omega_{I}\omega_{P_{2}}}}{4\epsilon_{0}v_{I}v_{P_{2}}\mathcal{L}^{(2)}}\int\mathrm{d}\boldsymbol{r}\Gamma_{ijkl}^{(3)}(\boldsymbol{r})\left[{\mathsf{d}}^{(2),i}_{S_{2}}(\boldsymbol{r}){\mathsf{d}}^{(2),j}_{S_{3}}(\boldsymbol{r})\right]^{*}{\mathsf{d}}^{(2),k}_{I}(\boldsymbol{r}){\mathsf{d}}^{(2),l}_{P_{2}}(\boldsymbol{r})e^{i(\kappa^{(2)}_{I}+\kappa^{(2)}_{P_{2}}-\kappa^{(2)}_{S_{3}}-\kappa^{(2)}_{S_{2}})\zeta}.$ (39)
Again, we note that to arrive to such expressions, full permutation symmetry of the $\Gamma^{(3)}$ tensor is assumed, which results from assuming that the indices of the second- and third-order nonlinear susceptibilities can be permuted freely [39]. For the materials of interest in this work, the contribution of the second-order nonlinear susceptibility to the $\Gamma^{(3)}$ tensor is negligible, thus we can write the $\Gamma^{(3)}$ tensor components as [39]:
$\Gamma^{(3)}_{ijkl}(\boldsymbol{r})=\frac{\chi^{(3)}_{ijkl}(\boldsymbol{r})}{\epsilon_{0}^{2}\varepsilon_{1}^{4}(\boldsymbol{r})}.$ (40)
@@
In identifying the effective nonlinear coefficient we neglect any frequency dependence on the dielectric constants, since we consider frequencies in the same range [39]. The spatial integral is restricted to the volume of the ring: We assume that the nonlinear interaction occurs only inside the ring, due to the field enhancement there compared to the waveguide, and since the nonlinear susceptibilities of the SiO2 cladding are orders of magnitude smaller than those of the material that constitutes the ring. We can relate the linear displacement fields to the electric fields
$\bm{\mathsf{d}}^{(\ell)}_{u}(\boldsymbol{r})=\epsilon_{0}\varepsilon_{1}(\boldsymbol{r})\bm{\mathsf{e}}^{(\ell)}_{u}(\boldsymbol{r}_{\perp},\zeta),$ (41)
where $\boldsymbol{r}_{\perp}$ is a vector in the cross-section perpendicular to the propagation direction of the field $\zeta$, and we can write
@@
$\begin{split}\gamma_{\mathrm{NL}}^{(\ell)}&=\frac{3\epsilon_{0}\sqrt{\omega_{{u_{3}}}\omega_{{u_{4}}}}}{4v_{{u_{3}}}v_{{u_{4}}}\mathcal{L}^{(\ell)}}\int\mathrm{d}\boldsymbol{r}_{\perp}\mathrm{d}\zeta\chi^{(3)}_{ijkl}(\boldsymbol{r})\left[{\mathsf{e}}^{(\ell),i}_{{u_{1}}}(\boldsymbol{r}_{\perp}){\mathsf{e}}^{(\ell),j}_{{u_{2}}}(\boldsymbol{r}_{\perp})\right]^{*}{\mathsf{e}}^{(\ell),k}_{{u_{3}}}(\boldsymbol{r}_{\perp}){\mathsf{e}}^{(\ell),l}_{{u_{4}}}(\boldsymbol{r}_{\perp})e^{i(\kappa^{(\ell)}_{{u_{3}}}+\kappa^{(\ell)}_{{u_{4}}}-\kappa^{(\ell)}_{{u_{2}}}-\kappa^{(\ell)}_{{u_{1}}})\zeta},\end{split}$ (42)
where the components of $\boldsymbol{{\mathsf{e}}}^{(\ell)}_{u}(\boldsymbol{r}_{\perp})$ can be extracted from an eigenmode solver for the specific structure and the integral over $\boldsymbol{r}_{\perp}$ can be evaluated numerically.
We have dropped the $\zeta$ dependence on the electric fields, writing only $\boldsymbol{\mathsf{e}}^{(\ell)}_{u}(\boldsymbol{r}_{\perp})$. This implicitly implies that the components of the fields $\boldsymbol{\mathsf{e}}^{(\ell)}_{u}(\boldsymbol{r}_{\perp})$ and of the tensor $\chi^{(3)}_{ijkl}$ are in the frame rotating with the light’s propagation along the rings: For each ring, we define this as the ring frame with coordinates xyzxyz (see Fig. 1a). In the ring frame, a vector $\boldsymbol{r}_{\perp}$ lies in the $xz$ plane, and $\zeta=\phi R_{\ell}$ goes along the $y$ axis. The integral over $\boldsymbol{r}_{\perp}$ ranges over the cross-section of the ring, and the one over $\zeta$ ranges from 0 to $\mathcal{L}^{(\ell)}$. For the interactions of interest, we consider processes with phase matched wavenumbers $\kappa^{(\ell)}_{{u_{3}}}+\kappa^{(\ell)}_{{u_{4}}}-\kappa^{(\ell)}_{{u_{2}}}-\kappa^{(\ell)}_{{u_{1}}}=0$; thus we can write
@@
$\begin{split}\gamma_{\mathrm{NL}}^{(\ell)}&=\frac{3\epsilon_{0}\sqrt{\omega_{{u_{3}}}\omega_{{u_{4}}}}}{4v_{{u_{3}}}v_{{u_{4}}}\mathcal{L}^{(\ell)}}\int\mathrm{d}\boldsymbol{r}_{\perp}\mathrm{d}\zeta\chi^{(3)}_{ijkl}(\boldsymbol{r})\left[{\mathsf{e}}^{(\ell),i}_{{u_{1}}}(\boldsymbol{r}_{\perp}){\mathsf{e}}^{(\ell),j}_{{u_{2}}}(\boldsymbol{r}_{\perp})\right]^{*}{\mathsf{e}}^{(\ell),k}_{{u_{3}}}(\boldsymbol{r}_{\perp}){\mathsf{e}}^{(\ell),l}_{{u_{4}}}(\boldsymbol{r}_{\perp}).\end{split}$ (43)
The III-V semiconductors of interest, such as AlGaAs, belong to the $\bar{4}3m$ point group and have cubic symmetry [6]. In general, for crystals in this point group the third-order nonlinear susceptibility $\chi^{(3)}$ tensor has 21 nonzero components, of which 4 are independent. We introduce the lab frame (axes $x^{\prime}y^{\prime}z^{\prime}$), which coincides with the principal axes of the crystal, and the nonzero tensor components of the $\chi^{(3)}$ tensor are ([6]):
$\chi^{(3)}_{x^{\prime}x^{\prime}x^{\prime}x^{\prime}}=\chi^{(3)}_{y^{\prime}y^{\prime}y^{\prime}y^{\prime}}=\chi^{(3)}_{z^{\prime}z^{\prime}z^{\prime}z^{\prime}}$, (44)
$\chi^{(3)}_{x^{\prime}x^{\prime}y^{\prime}y^{\prime}}=\chi^{(3)}_{y^{\prime}y^{\prime}x^{\prime}x^{\prime}}=\chi^{(3)}_{y^{\prime}y^{\prime}z^{\prime}z^{\prime}}=\chi^{(3)}_{z^{\prime}z^{\prime}y^{\prime}y^{\prime}}=\chi^{(3)}_{z^{\prime}z^{\prime}x^{\prime}x^{\prime}}=\chi^{(3)}_{x^{\prime}x^{\prime}z^{\prime}z^{\prime}}$, (45)
$\chi^{(3)}_{x^{\prime}y^{\prime}x^{\prime}y^{\prime}}=\chi^{(3)}_{y^{\prime}x^{\prime}y^{\prime}x^{\prime}}=\chi^{(3)}_{y^{\prime}z^{\prime}y^{\prime}z^{\prime}}=\chi^{(3)}_{z^{\prime}y^{\prime}z^{\prime}y^{\prime}}=\chi^{(3)}_{z^{\prime}x^{\prime}z^{\prime}x^{\prime}}=\chi^{(3)}_{x^{\prime}z^{\prime}x^{\prime}z^{\prime}}$, (46)
$\chi^{(3)}_{x^{\prime}y^{\prime}y^{\prime}x^{\prime}}=\chi^{(3)}_{y^{\prime}x^{\prime}x^{\prime}y^{\prime}}=\chi^{(3)}_{y^{\prime}z^{\prime}z^{\prime}y^{\prime}}=\chi^{(3)}_{z^{\prime}y^{\prime}y^{\prime}z^{\prime}}=\chi^{(3)}_{z^{\prime}x^{\prime}x^{\prime}z^{\prime}}=\chi^{(3)}_{x^{\prime}z^{\prime}z^{\prime}x^{\prime}}$. (47)
The components can be written as ([46, 24])
$\chi^{(3)}_{i^{\prime}j^{\prime}k^{\prime}l^{\prime}}=\chi^{(3)}_{x^{\prime}x^{\prime}y^{\prime}y^{\prime}}\delta_{i^{\prime}j^{\prime}}\delta_{k^{\prime}l^{\prime}}+\chi^{(3)}_{x^{\prime}y^{\prime}x^{\prime}y^{\prime}}\delta_{i^{\prime}k^{\prime}}\delta_{j^{\prime}l^{\prime}}+\chi^{(3)}_{x^{\prime}y^{\prime}y^{\prime}x^{\prime}}\delta_{i^{\prime}l^{\prime}}\delta_{j^{\prime}k^{\prime}}+(\chi^{(3)}_{x^{\prime}x^{\prime}x^{\prime}x^{\prime}}-\chi^{(3)}_{x^{\prime}x^{\prime}y^{\prime}y^{\prime}}-\chi^{(3)}_{x^{\prime}y^{\prime}x^{\prime}y^{\prime}}-\chi^{(3)}_{x^{\prime}y^{\prime}y^{\prime}x^{\prime}})\delta_{i^{\prime}j^{\prime}k^{\prime}l^{\prime}},$ (48)
where $\delta_{ab}$ is the usual Kronecker delta function, and $\delta_{abcd}$ is equal to unity if $a=b=c=d$, and zero otherwise.
In arriving at expressions for the nonlinear parameters such as Eq. (42), permutation symmetry of the indices of the $\Gamma^{(3)}$ tensor was assumed, and for AlGaAs this is analogous to assuming Kleinman symmetry. Under this approximation, Eqs. (45) to (47) are all equal, which implies that the third-order nonlinear susceptibility tensor has only two independent components: $\chi^{(3)}_{x^{\prime}x^{\prime}x^{\prime}x^{\prime}}$ and $\chi^{(3)}_{x^{\prime}x^{\prime}y^{\prime}y^{\prime}}$. Thus, under Kleinman symmetry, we can write the tensor components as
$\chi^{(3)}_{i^{\prime}j^{\prime}k^{\prime}l^{\prime}}=\chi^{(3)}_{x^{\prime}x^{\prime}y^{\prime}y^{\prime}}(\delta_{i^{\prime}j^{\prime}}\delta_{k^{\prime}l^{\prime}}+\delta_{i^{\prime}k^{\prime}}\delta_{j^{\prime}l^{\prime}}+\delta_{i^{\prime}l^{\prime}}\delta_{j^{\prime}k^{\prime}})+(\chi^{(3)}_{x^{\prime}x^{\prime}x^{\prime}x^{\prime}}-3\chi^{(3)}_{x^{\prime}x^{\prime}y^{\prime}y^{\prime}})\delta_{i^{\prime}j^{\prime}k^{\prime}l^{\prime}}.$ (49)
In this framework, the third-order nonlinear optical interactions can be described by two independent components of the nonlinear susceptibility: $\chi^{(3)}_{x^{\prime}x^{\prime}x^{\prime}x^{\prime}}$ and $\chi^{(3)}_{x^{\prime}x^{\prime}y^{\prime}y^{\prime}}$. As mentioned in Johnson [24], assuming Kleinman symmetry for frequencies close to the half band gap (such as those considered in this work) often leads to underestimating the nonlinear interaction strength; thus we anticipate that our calculated results (e.g., generation rates) to be slightly lower than those obtainable from experiments in fabricated samples.
From the form that we have for the nonlinear parameters in Eq. (42), to contract the indices using the electric field mode profiles in the ring frame we need the tensor components of $\chi^{(3)}$ in the ring frame. For a vector $\boldsymbol{v}$, with components $v_{i}$ and $v_{i^{\prime}}$ in the ring and lab frames, respectively, we can relate its components according to
$\begin{split}v_{i}&=\boldsymbol{M}_{\phi}^{i{i^{\prime}}}v_{i^{\prime}},\\ v_{i^{\prime}}&=\boldsymbol{M}_{\phi}^{ii^{\prime}}v_{i},\end{split}$ (50)
where
$\boldsymbol{M}_{\phi}=\begin{bmatrix}\cos\phi&\sin\phi&0\\ -\sin\phi&\cos\phi&0\\ 0&0&1\end{bmatrix}$. (51)
It follows that for a rank-$n$ tensor, we can identify its components in the ring frame with
$\Upsilon^{(n)}_{i_{1},\dots,i_{n}}=\boldsymbol{M}_{\phi}^{i_{1}{i_{1}^{\prime}}}\dots\boldsymbol{M}_{\phi}^{i_{n}{i_{n}^{\prime}}}\Upsilon^{(n)}_{i^{\prime}_{1},\dots,i^{\prime}_{n}}$. (52)
To simplify the discussion, we specialize to the case considered in the text where all four fields are TE00 polarized. This implies that to good approximation the main component of the fields are $\mathsf{e}^{x}_{u}$, and all of the terms involving $\mathsf{e}^{y}_{u}$ and $\mathsf{e}^{z}_{u}$ are near zero. In this case, only the contraction of the fields in the ring frame with the $\chi^{(3)}_{xxxx}$ tensor component contribute to the integral for the nonlinear parameters from Eq. (43). To good approximation we can write
$\begin{split}\chi^{(3)}_{ijkl}(\boldsymbol{r})\left[{\mathsf{e}}^{i}_{{u_{1}}}{\mathsf{e}}^{j}_{{u_{2}}}\right]^{*}{\mathsf{e}}^{k}_{{u_{3}}}{\mathsf{e}}^{l}_{{u_{4}}}\approx\left[{\mathsf{e}}^{x}_{{u_{1}}}{\mathsf{e}}^{x}_{{u_{2}}}\right]^{*}{\mathsf{e}}^{x}_{{u_{3}}}{\mathsf{e}}^{x}_{{u_{4}}}\times\left(A-\frac{F}{2}\sin^{2}(2\phi)\right),\end{split}$ (53)
where we defined
$A$ $\equiv\chi^{(3)}_{x^{\prime}x^{\prime}x^{\prime}x^{\prime}}$, (54)
$B$ $\equiv\chi^{(3)}_{x^{\prime}x^{\prime}y^{\prime}y^{\prime}}$, (55)
$F$ $\equiv A-3B.$ (56)
The effect of the anisotropy on the nonlinear parameter is described by the term with $F$, where if $F>0$ the anisotropy would reduce the nonlinear parameter and if $F<0$ it would increase the nonlinear parameter, since $\int\mathrm{d}\zeta\sin^{2}(2\phi)>0$.
The magnitude of $F$, and hence the effect of the anisotropy, depends on the material considered. Experimental work and modelling (without assuming Kleinman symmetry) was done for AlGaAs with 18% aluminum content [23, 22, 24], and we assume that the AlGaAs we consider in this manuscript (30% aluminum content) has similar properties. When the frequencies of the generated light are close to each other, as is the case in this work, to good approximation the tensor components $\chi^{(3)}_{x^{\prime}y^{\prime}x^{\prime}y^{\prime}}$ and $\chi^{(3)}_{x^{\prime}y^{\prime}x^{\prime}y^{\prime}}$ can be set to be equal, in which case one can introduce [23, 22, 24] the anisotropy parameter $\sigma=(\chi^{(3)}_{x^{\prime}x^{\prime}x^{\prime}x^{\prime}}-\chi^{(3)}_{x^{\prime}x^{\prime}y^{\prime}y^{\prime}}-2\chi^{(3)}_{x^{\prime}y^{\prime}x^{\prime}y^{\prime}})/(2\chi^{(3)}_{x^{\prime}x^{\prime}x^{\prime}x^{\prime}})$, which characterizes the effect of the anisotropy similar to the $F$ we introduced in Eq. (56). In the cited studies, the values of $\sigma$ are negative (and hence $F<0$), implying that the anisotropy of the material effectively increases the nonlinear parameters. For this manuscript, to ensure that we do not overestimate the nonlinear parameters, we use $F=0$; we expect the nonlinear parameters to be slightly larger than those we compute. A more thorough investigation of the change of the magnitudes of the nonlinear parameters from the anisotropy could be done, although significant enhancements of the nonlinear parameters are not expected.
Thus, for the four-wave mixing processes in AlGaAs that we consider in this manuscript, we can compute the nonlinear coefficients to good approximation with
$\begin{split}\gamma_{\mathrm{NL}}^{(\ell)}&\approx\frac{3\chi^{(3)}_{x^{\prime}x^{\prime}x^{\prime}x^{\prime}}\epsilon_{0}\sqrt{\omega_{{u_{3}}}\omega_{{u_{4}}}}}{4v_{{u_{3}}}v_{{u_{4}}}}\int\mathrm{d}\boldsymbol{r}_{\perp}\left[{\mathsf{e}}^{(\ell),x}_{{u_{1}}}(\boldsymbol{r}_{\perp}){\mathsf{e}}^{(\ell),x}_{{u_{2}}}(\boldsymbol{r}_{\perp})\right]^{*}{\mathsf{e}}^{(\ell),x}_{{u_{3}}}(\boldsymbol{r}_{\perp}){\mathsf{e}}^{(\ell),x}_{{u_{4}}}(\boldsymbol{r}_{\perp}).\end{split}$ (57)
Appendix D Asymptotic fields for dual rings coupled to a waveguide
$\boldsymbol{D}(\boldsymbol{r})=\sum_{u}\boldsymbol{D}_{u}^{\mathrm{in/out}}(\boldsymbol{r})+\mathrm{H.c.},$ (58)
with
$\boldsymbol{D}_{u}^{\mathrm{in/out}}(\boldsymbol{r})=\sum_{\lambda}\int\mathrm{d}k\boldsymbol{\mathcal{D}}_{uk}^{\mathrm{in/out},\lambda}(\boldsymbol{r})a_{u}^{\mathrm{in/out},\lambda}(k)+\mathrm{H.c..}$ (59)
We provide the fields inside the microring resonators for each case in Fig. 4, obtained from the Heisenberg equations for the specific asymptotic state [5]. Using $\boldsymbol{r}_{\perp}$ to denote a vector along the cross-section of the waveguide constructing the ring, and $\zeta$ to identify the azimuthal coordinate along the circumference of the ring (see Fig. 1a), the asymptotic-in fields for the actual waveguide (Fig. 4a) are
$\boldsymbol{\mathcal{D}}_{uk}^{\mathrm{in,ac}}(\boldsymbol{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\boldsymbol{\mathsf{d}}^{(1)}_{u}(\boldsymbol{r}_{\perp})F_{u-}^{\mathrm{ac}}(k)e^{-i(k-K_{u})a}e^{i\kappa^{(1)}_{u}\zeta},$
$[\boldsymbol{r}\in\,\mathrm{ring\,1}]$ (60)
$\boldsymbol{\mathcal{D}}_{uk}^{\mathrm{in,ac}}(\boldsymbol{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\boldsymbol{\mathsf{d}}^{(2)}_{u}(\boldsymbol{r}_{\perp})F_{u-}^{\mathrm{ac}(2)}(k)\left(1+\frac{i\gamma_{u}^{\mathrm{ac}(1)}}{v_{u}^{\mathrm{ac}}}\sqrt{\mathcal{L}^{(1)}}F_{u-}^{\mathrm{ac}(1)}(k)\right)e^{i(k-K_{u})a}e^{i\kappa^{(2)}_{u}\zeta},$
$[\boldsymbol{r}\in\,\mathrm{ring\,2}]$ (61)
the asymptotic-out fields for the actual waveguide (Fig. 4b) are
$\boldsymbol{\mathcal{D}}_{uk}^{\mathrm{out,ac}}(\boldsymbol{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\boldsymbol{\mathsf{d}}^{(1)}_{u}(\boldsymbol{r}_{\perp})F_{u+}^{\mathrm{ac}(1)}(k)\left(1-\frac{i\gamma_{u}^{\mathrm{ac}(2)}}{v_{u}^{\mathrm{ac}}}\sqrt{\mathcal{L}^{(2)}}F_{u+}^{\mathrm{ac}(2)}(k)\right)e^{-i(k-K_{u})a}e^{i\kappa^{(1)}_{u}\zeta},$
$[\boldsymbol{r}\in\,\mathrm{ring\,1}]$ (62)
$\boldsymbol{\mathcal{D}}_{uk}^{\mathrm{out,ac}}(\boldsymbol{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\boldsymbol{\mathsf{d}}^{(2)}_{u}(\boldsymbol{r}_{\perp})F_{u+}^{\mathrm{ac}(2)}(k)e^{i(k-K_{u})a}e^{i\kappa^{(2)}_{u}\zeta},$
$[\boldsymbol{r}\in\,\mathrm{ring\,2}]$ (63)
$\boldsymbol{\mathcal{D}}_{uk}^{\mathrm{out,ph}(1)}(\boldsymbol{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\boldsymbol{\mathsf{d}}^{(1)}_{u}(\boldsymbol{r}_{\perp})F_{u+}^{\mathrm{ph}(1)}(k)e^{-i(k-K_{u})a}e^{i\kappa^{(1)}_{u}\zeta},$
$[\boldsymbol{r}\in\,\mathrm{ring\,1}]$ (64)
$\boldsymbol{\mathcal{D}}_{uk}^{\mathrm{out,ph}(1)}(\boldsymbol{r}_{\perp})=0,$
$[\boldsymbol{r}\in\,\mathrm{ring\,2}]$ (65)
$\boldsymbol{\mathcal{D}}_{uk}^{\mathrm{out,ph}(2)}(\boldsymbol{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\boldsymbol{\mathsf{d}}^{(1)}_{u}(\boldsymbol{r}_{\perp})F_{u+}^{\mathrm{ac}(1)}(k)\left(-\frac{i\gamma_{u}^{\mathrm{ac}(2)}}{v_{u}^{\mathrm{ac}}}\sqrt{\mathcal{L}^{(2)}}F_{u+}^{\mathrm{ph}(2)}(k)\right)e^{-i(k-K_{u})a}e^{i\kappa^{(1)}_{u}\zeta}$
$[\boldsymbol{r}\in\,\mathrm{ring\,1}]$ (66)
$\boldsymbol{\mathcal{D}}_{uk}^{\mathrm{out,ph}(2)}(\boldsymbol{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\boldsymbol{\mathsf{d}}^{(2)}_{u}(\boldsymbol{r}_{\perp})F_{u+}^{\mathrm{ph}(2)}(k)e^{i(k-K_{u})a}e^{i\kappa^{(2)}_{u}\zeta},$
$[\boldsymbol{r}\in\,\mathrm{ring\,2}]$ (67)
where $\boldsymbol{\mathsf{d}}^{(\ell)}_{u}(\boldsymbol{r})$ are the spatial mode profiles in the $\ell^{\mathrm{th}}$ ring ($\ell\in\{1,2\}$) associated with the frequency range $u$ [39]. We have also introduced a field enhancement factor
The spatial modes are normalized according to
$\begin{split}\int\frac{v_{p}(\boldsymbol{r}_{\perp};\omega_{u})}{v_{g}(\boldsymbol{r}_{\perp};\omega_{u})}\frac{\bm{\mathsf{d}}_{u}^{(\ell)*}(\boldsymbol{r}_{\perp})\cdot\bm{\mathsf{d}}_{u}^{(\ell)}(\boldsymbol{r}_{\perp})}{\epsilon_{0}\varepsilon_{1}(\boldsymbol{r}_{\perp};\omega_{u})}\mathrm{d}\boldsymbol{r}_{\perp}=1,\end{split}$ (69)